OpenStudy (anonymous):

Help? :c Use mathematical induction to prove the statement is true for all positive integers n. The integer n^3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

plug in (x+1) into the formula n^3 + 2n what do you get ?

OpenStudy (anonymous):

\[x^3+3x^2+6x+4\]

OpenStudy (phi):

that does not look right. can you show your steps ?

OpenStudy (anonymous):

My bad. \[x^3+3x^3+5x+3\]

OpenStudy (phi):

ok, now with that accomplished, we can do the proof by induction. 1) show that \[ n^3 +2n \text{ is evenly divisible by 3}\] is true for n=1

OpenStudy (anonymous):

1^3+2(1)=1+2=3

OpenStudy (phi):

and 3 is evenly divisible by 3 next step 2) assume \[ n^3 +2n \text{ is evenly divisible by 3} \] for all numbers x ≤ n

OpenStudy (phi):

we use step 2 this way: we let x= n+1, and use that value in the formula. As you figured out, we get \[ n^3+3n^3+5n+3 \] (sorry about switching the n's and x's) we can rewrite your result using 5n= 3n + 2n \[ n^3+2n + 3n^3+3n+3 \]

OpenStudy (phi):

notice that the first 2 terms \[ n^3 +2n\] is assumed to be divisible by 3. (In other words, according to step 2, we assume the formula works for all numbers ≤ n.) also, we see the last 3 terms \[ 3n^3 +3n+3= 3(n^3+n+1) \] is evenly divisible by 3 therefore the formula worked for x= n+1

OpenStudy (phi):

in summary 1) formula is divisible by 3 for n=1 2) formula is divisible by 3 for n= n+1 therefore the formula is true for all n≥1

OpenStudy (anonymous):

How did we prove that n^3+2n was?

OpenStudy (phi):

We don't prove that n^3+2n is evenly divisible by 3. We ASSUME it is true for some n what we prove is that IF n^3+2n is divisible by 3, then it is also true for x= n+1 (which we did) the idea is that if n^3+2n is divisible by 3, and it is also true for n= n+1 (no matter what n is), then the formula works for n+2, n+3, n+4, and so on....

OpenStudy (phi):

by step 1, we know the formula works for n=1 by step 2, if the formula works for n=1, it will work for n=1+1 = 2 if the formula works for n=2 (it does) then by step 2, it works for n=2+1=3 if the formula works for n=3, then by step 2, it works for n=4 and we keep going forever.... because the formula works for n+1, we can prove it works for each number in turn, forever...

OpenStudy (anonymous):

Oh, just got it. Had to stare at it for a while. Thanks so much! :DDD God bless you!

OpenStudy (phi):

Here is more background http://www.khanacademy.org/math/precalculus/seq_induction/proof_by_induction/v/proof-by-induction It is a tricky idea.